# Lensing and wave optics in the strong field of a black hole

### v2

@@ I Introduction
Gravitational lensing offers a powerful tool to study wave propagation in curved spacetime. While extensively observed for electromagnetic waves, the lensing of gravitational waves (GWs) remains undetected. However, both upgraded current-generation detectors [1, 2, 3] and future observatories are expected to routinely detect lensed GW events [4, 5]. GWs undergo lensing in a similar way to light but with key differences: (i) their longer wavelengths allow us to probe novel gravitational lensing regimes [6], and (ii) their spin-2 nature leads to two polarization states, $+$ and $\times$, which can mix during propagation through strongly curved spacetimes, resulting in distinctive observational signatures  [7, 8, 9, 10].
@@
Wave optics effects in GW lensing have been extensively studied under the assumption that the lens produces only a weak gravitational potential. This approximation is valid for large-scale structures such as dark matter halos, galaxies, and galaxy clusters [15]. However, it becomes less accurate when the lens is a black hole (BH), where strong-field effects play a crucial role. In the Newtonian regime, an analytic solution exists under the point-mass lens (PL) approximation [16, 11], which is particularly useful to study GW lensing [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 12, 13, 35, 36, 37, 38]. However, this approximation neglects the strong gravitational potential near a BH and finite-size effects. Recent studies have taken steps toward addressing this limitation: numerical analyses of GW lensing by a non-rotating BH have revealed interesting magnification effects [39], while deep wave optics analyses have explored the problem in the long-wavelength regime [40]. Later studies have extended this to the case of spinning black holes [41, 42].
@@
In this work, we investigate GW lensing in the wave optics regime, comparing the PL approximation to the full strong-field regime. Our results extend those of [40] by going beyond the deep wave optics limit and improve upon [39] by introducing a more efficient formalism that allows rapid evaluation of lensed GW signals for arbitrary source-lens-observer configurations. We construct an amplification factor that incorporates strong-field effects and explicitly accounts for the mixing between the $+$ and $\times$ GW polarizations. Finally, we demonstrate that while the PL approximation captures some qualitative features of BH-lensed GWs, significant differences arise due to polarization mixing and BH-induced absorption.
This paper is structured as follows. In Sec. III we review GW lensing in the weak field regime, and the PL approximation. In Sec. II we examine GW lensing in the strong field of a non-rotating BH. Our main results are presented in Sec. IV, followed by a discussion of future research directions in Sec. V. Throughout this work, we adopt the Planck 2018 cosmology [55], and unless otherwise specified, we work with geometric units, $G=1=c$.
@@ II GW lensing in weak fields
Figure 1: Diagram exemplifying the distinct regimes of GW lensing. We denote $M_{L}$ the lens mass, $\mathcal{R}$, its characteristic size and, $\lambda$ the wavelength of the incoming wave. In this work we are primarily concerned with the strong gravity, wave optics regime (top-left), particularly in the application of BH perturbation theory to explore the phenomenology of GW distortions in this part of parameter space. We will also make an explicit connection to lensing characteristics in the weak gravity, wave optics regime (top-right).
@@
Moreover, it is customary to consider the lens to be thin, relative to the characteristic distances of the problem. This allows us to approximate the lens potential $\Phi=\delta(r-r_{l})\phi(\vec{x})$, $\vec{x}$ being the image position on the lens plane, and $r_{l}$ the distance from the observer to the lens [56]. The propagated of the GWs is only affected by a two dimensional projection $\phi(\vec{x})$ of the potential.
@@
×$\times\exp[i\varpi\Bigl{(}\frac{x^{2}}{2}-\frac{\phi(x)}{2M}\Bigr{)}\Bigr{]}\,.$
$F(\varpi,y)=$ $\exp\bigg{[}\dfrac{\pi\varpi}{4}+i\dfrac{\varpi}{2}\bigg{\{}\ln\bigg{(}\dfrac{\varpi}{2}\bigg{)}-2\phi_{m}(y)\bigg{\}}\bigg{]}$
×$\times\Gamma\bigg{(}1-\dfrac{i}{2}\varpi\bigg{)}{{}_{1}}F_{1}\bigg{(}\dfrac{i}{2}\varpi,1;\dfrac{i}{2}\varpi y^{2}\bigg{)}\,,$ (6)
@@
The description of GW lensing by a PL is particularly simple. However, much information is lost through the several approximations being made here. For instance, despite GWs being a spin $2$ wave, both polarizations are lensed equally. This is seen explicitly in Eq. (II), since the amplification factor $F$ does not carry information about the helicity of the GWs. Additionally, the finite size of the BH is neglected, and as a consequence there is no way to account for absorption of GWs by the BH itself. All that Eq. (II) captures is the diffraction effects of a wave scattered by a point-like potential. As we will see in following sections, the effects neglected here can be captured by treating the strong field of a BH accurately, at the expense of some technical difficulties.
@@ III.1 GWs Emitted by the source
We first describe the GWs emitted by the source in its own reference frame. The source consists of two merging black holes, with a merger timescale short enough that the relative motion between the source and the lens can be neglected [62, 63, 64]. In the transverse traceless (TT) gauge, the GWs emitted by the source can be written as
@@
where $h_{+(\times)}$ are the strain amplitudes of the $+$ ($\times$) polarizations, in Cartesian $(t,x,y,z)$ coordinates. In the frequency domain, the strain amplitudes take the form
@@
where $\omega=2\pi f(1+z)$ is the angular frequency at the source location, written in terms of the detector frequency $f$ measured at the observer, and $\mathfrak{h}^{LM}_{+/\times}(f)$ are the (real valued) strain amplitudes for each polarization. The angles $\Omega_{\text{LS}}=\{\theta_{\text{LS}},\phi_{\text{LS}}\}$, describe the lens’ position relative to the observer. The $+$ and $\times$ polarizations are not helicity eigenstates, as they lack a well defined spin weight. However helicity eigenstates can be obtained via
@@ III.2 BH Scattering
with $M$ the BH mass, and $d\Omega^{2}$ the line element on the unit sphere. The linearized Einstein equations reduce to two wave equations for each spherical harmonic, the Regge-Wheeler (RW) and Zerilli equations [53, 54]
@@
where =$\bullet=\{\mathrm{odd},\mathrm{even}\}$. The respective potentials are given by
@@
where the wave is purely in-going at the horizon, $\psi\sim e^{-i\omega r_{\star}}$ as $r\to 2M$ ($r_{\star}\to-\infty$). We define the reflectivity of the BH as
@@
We implemented two methods to compute reflectivity coefficients across frequencies: one leveraging the Black Hole Perturbation Toolkit implementation of the Mano-Suzuki-Takasugi (MST) method [71, 72], and another via direct RW equation integration. Both methods agree to sub-percent accuracy up to $\ell=30$, while the MST method enables reliable computations up to $\ell_{\rm max}=70$ for $\omega M\in[0.01,14]$. The reflectivity behavior is shown in Fig. 3. Notice that the reflectivity in the even parity sector is equal in magnitude to the odd sector (as required by isospectrality), and with a phase shift which can be computed through
@@
The well defined helicity modes of the GW strain $h_{(\pm 2)}$ can be decomposed in spin weighted spherical harmonics centered around the lens
@@
×$\quad\times\Bigl{[}(-1)^{\ell}e^{-i\omega r_{\star}}-e^{i\omega r_{\star}}\Bigr{]}+\mathrm{c.m.}\,,$
@@
This solution corresponds to the incoming spin-2 wave impinging the BH and passing through the system unaffected. We account for the strong field of the BH by rescaling the out-going field with the BH reflectivity factors. Doing so, and subtracting the incoming waveform (which we will recover at a later point), we find
@@
×$\quad\times\Bigl{[}1+(-1)^{\ell}\mathcal{R}^{\rm even}_{\ell m}\Bigr{]}e^{i\omega r_{\star}}+\mathrm{c.m.}\,,$
@@
$\tilde{H}$ =(±${}^{(\pm 2)}=\mathcal{F}^{\pm}H^{(\mp 2)}/\omega\,,$ (28)
@@
$f$ =${}^{\pm}_{\ell m}=\Bigl{[}1+(-1)^{\ell}\mathcal{R}^{\rm even}_{\ell m}\Bigr{]}\mp\frac{m}{|m|}\Bigl{[}1+(-1)^{\ell}\mathcal{R}^{\rm odd}_{\ell m}\Bigr{]}\,,$
@@
For each $\ell$ mode, there are two contributions, coming from $m=\pm s$, with $s=\pm 2$ in this case. At large $\ell$, the even and odd reflectivities become nearly identical, i.e., $\mathcal{R}_{\ell m}^{\rm even}-\mathcal{R}_{\ell m}^{\rm odd}\xrightarrow{\ell\to\infty}0$. As a result, the partial sums for $m=s$, which depend on the difference between these reflectivities, exhibit rapid convergence. In contrast, for $m=-s$ the partial sums do not converge. This is a known phenomenon in the multipolar expansion of the scattering of plane waves [75, 40], and convergence is only achieved within some finite radius of the BH for a given value of $\ell_{\text{max}}$ [76]. Nevertheless, while the sum does not converge in the conventional sense, it possesses a well-defined Cesàro sum  [^3] The Cesàro sum, in its basic form, is obtained by averaging the sequence of partial sums and taking the limit. More details are given in the Appendix A., which we have verified numerically. In the Supplemental Material, we provide extensive numerical evidence of the sum’s effective convergence, as well as the errors introduced when truncating at finite $\ell$. We note that achieving convergence requires summing over a larger range of $\ell$ modes when the observer-lens angle $\theta_{L}$ is small. This imposes a practical limitation on probing smaller impact parameters, where richer strong-field effects will be observed.
@@ III.3 Waveform at the observer’s location
where $r_{\rm OS},\Omega_{\rm OS}$ are, respectively, the distance and the solid angle connecting the source and the observer’s locations. We note that this formulation assumes no interaction between the direct component of the emitted gravitational wave and the potential of the lens, serving as the primary simplifying assumption of this work. A more sophisticated treatment of the direct piece must be required in order to recover the geometric optics results of the PL model in the high-frequency limit. We focus on cases where the lens and source are relatively close compared to the observer, $r_{\rm LS}\ll r_{\rm OS}\sim r_{\rm OL}$. Under this approximation $\Omega_{\rm OS}\simeq\Omega_{\rm LS}$ (with equality holding in the case where all source, lens and observer are aligned). This allows us to express the total waveform compactly as
@@
The total waveform is $h=h_{+}\mathcal{A}_{+}+h_{\times}\mathcal{A}_{\times}$, where $\mathcal{A}_{+/\times}$ are the detector’s antenna response functions of the $+$ and $\times$ polarizations. Including the effects of lensing, this becomes
@@
Figure 4: Absolute value of the amplification factor relative to the non-magnified case, for the polarization preserving contribution, $|\mathcal{F}-1|^{2}$ (blue), and of the total polarization-mixing amplification $|\mathcal{G}|^{2}$ (green), as a function of the dimensionless frequency $\omega M$. The lens is located at $r_{\rm SL}=100M$, and the source has angular coordinates $(\theta_{\rm OS},\phi_{\rm OS})=(\pi/6,\pi/6)$ with respect to the observer. In red, we give $|\mathcal{F}-1|^{2}$ for the case of a PL, corresponding to an impact parameter $y=2.69$. The dimensionless frequency $\varpi$ is related to $\omega M$ via Eq. (2).
In Fig. 4, we additionally show the square of $\mathcal{F}-1$ computed using the point mass lens (PL) approximation, for a situation matching the strong field calculation. We emphasize two main differences: (i) the strong field calculation outputs larger magnifications, relative to the PL approximation (for the polarization mixing contribution, due to our choice of the angle $\phi$), and (ii) at high frequencies, the amplification factor $|\mathcal{F}|\to 1$ and $\mathcal{G}\to 0$, since absorption by the BH becomes important. This effect is unaccounted for in the PL approximation. Nevertheless, it is remarkable that a simple approximation such as the PL can reproduce accurately several of the features observed in the complete, strong field calculation.
@@ IV Waveform comparison
Once we have constructed the amplification factors, we can directly compute the lensed waveform and analyze its most prominent features. Our amplification factor can be evaluated efficiently across a range of $\omega M$, and thus we capture key lensing effects in the strong gravitational regime throughout the entire inspiral-merger-ringdown. We will examine both the strong-field regime and the weak-field, PL approximation. For clarity, we will focus on a specific gravitational waveform resulting from the merger of spinless, equal mass BHs. This allows us to isolate the main contributions from strong-field lensing compared to the weak-field approximation. We will explore the lensed waveform in both the frequency and time domains.

### v3

@@ I Introduction
Gravitational lensing offers a powerful tool to study wave propagation in curved spacetime. While extensively observed for electromagnetic waves, the lensing of gravitational waves (GWs) remains undetected. However, both upgraded current-generation detectors [1, 2, 3] and future observatories are expected to routinely detect lensed GW events [4, 5]. GWs undergo lensing in a similar way to light but with key differences: (i) their longer wavelengths allow us to probe novel gravitational lensing regimes [6], and (ii) their spin-2 nature leads to two polarization states, ++ and $\times$, which can mix during propagation through strongly curved spacetimes, resulting in distinctive observational signatures  [7, 8, 9, 10].
@@
Wave optics effects in GW lensing have been extensively studied under the assumption that the lens produces only a weak gravitational potential. This approximation is valid for large-scale structures such as dark matter halos, galaxies, and galaxy clusters [15]. However, it becomes less accurate when the lens is a black hole (BH), where strong-field effects play a crucial role. In the Newtonian regime, an analytic solution exists under the point-mass lens (PL) approximation [16, 11], which is particularly useful to study GW lensing [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 12, 13, 35, 36, 37, 38]. However, this approximation neglects the strong gravitational potential near a BH and finite-size effects. Recent studies have taken steps toward addressing this limitation: numerical analyses of GW lensing by a nonrotating BH have revealed interesting magnification effects [39], while deep wave optics analyses have explored the problem in the long-wavelength regime [40]. Later studies have extended this to the case of spinning black holes [41, 42].
@@
In this work, we investigate GW lensing in the wave optics regime, comparing the PL approximation to the full strong-field regime. Our results extend those of [40] by going beyond the deep wave optics limit and improve upon [39] by introducing a more efficient formalism that allows rapid evaluation of lensed GW signals for arbitrary source-lens-observer configurations. We construct an amplification factor that incorporates strong-field effects and explicitly accounts for the mixing between the ++ and $\times$ GW polarizations. Finally, we demonstrate that while the PL approximation captures some qualitative features of BH-lensed GWs, significant differences arise due to polarization mixing and BH-induced absorption.
This paper is structured as follows. In Sec. III we review GW lensing in the weak field regime, and the PL approximation. In Sec. II we examine GW lensing in the strong field of a nonrotating BH. Our main results are presented in Sec. IV, followed by a discussion of future research directions in Sec. V. Throughout this work, we adopt the Planck 2018 cosmology [55], and unless otherwise specified, we work with geometric units, $G=1=c$.
@@ II GW lensing in weak fields
Figure 1: Diagram exemplifying the distinct regimes of GW lensing. We denote $M_{L}$ the lens mass, $\mathcal{R}$, its characteristic size and, $\lambda$ the wavelength of the incoming wave. In this work we are primarily concerned with the strong gravity, wave optics regime (topleft), particularly in the application of BH perturbation theory to explore the phenomenology of GW distortions in this part of parameter space. We will also make an explicit connection to lensing characteristics in the weak gravity, wave optics regime (topright).
@@
Moreover, it is customary to consider the lens to be thin, relative to the characteristic distances of the problem. This allows us to approximate the lens potential $\Phi=\delta(r-r_{l})\phi(\vec{x})$, $\vec{x}$ being the image position on the lens plane, and $r_{l}$ the distance from the observer to the lens [56]. The propagation of the GWs is only affected by a two-dimensional projection $\phi(\vec{x})$ of the potential.
@@
$F(\varpi,y)=$ $-i\varpi e^{i\varpi y^{2}/2}\int_{0}^{\infty}dxxJ_{0}(\varpi xy)$ (5)
$\times\exp[i\varpi\Bigl{(}\frac{x^{2}}{2}-\frac{\phi(x)}{2M}\Bigr{)}\Bigr{]}\,.$
@@
$\times\Gamma\bigg{(}1-\dfrac{i}{2}\varpi\bigg{)}{{}_{1}}F_{1}\bigg{(}\dfrac{i}{2}\varpi,1;\dfrac{i}{2}\varpi y^{2}\bigg{)}\,$, (6)
@@
The description of GW lensing by a PL is particularly simple. However, much information is lost through the several approximations being made here. For instance, despite GWs being a spin $2$ wave, both polarizations are lensed equally. This is seen explicitly in Eq. (II), since the amplification factor $F$ does not carry information about the helicity of the GWs. Additionally, the finite size of the BH is neglected, and as a consequence there is no way to account for absorption of GWs by the BH itself. All that Eq. (II) captures is the diffraction effects of a wave scattered by a pointlike potential. As we will see in following sections, the effects neglected here can be captured by treating the strong field of a BH accurately, at the expense of some technical difficulties.
@@ III.1 GWs emitted by the source
We first describe the GWs emitted by the source in its own reference frame. The source consists of two merging black holes, with a merger timescale short enough that the relative motion between the source and the lens can be neglected [62, 63, 64]. In the transverse traceless gauge, the GWs emitted by the source can be written as
@@
where $h_{+(\times)}$ are the strain amplitudes of the ++ ($\times$) polarizations, in Cartesian $(t,x,y,z)$ coordinates. In the frequency domain, the strain amplitudes take the form
@@
where $\omega=2\pi f(1+z)$ is the angular frequency at the source location, written in terms of the detector frequency $f$ measured at the observer, and $\mathfrak{h}^{LM}_{+/\times}(f)$ are the (real valued) strain amplitudes for each polarization. The angles $\Omega_{\text{LS}}=\{\theta_{\text{LS}},\phi_{\text{LS}}\}$, describe the lens’ position relative to the observer. The ++ and $\times$ polarizations are not helicity eigenstates, as they lack a well defined spin weight. However helicity eigenstates can be obtained via
@@ III.2 BH scattering
with $M$ the BH mass, and $d\Omega^{2}$ the line element on the unit sphere. The linearized Einstein equations reduce to two wave equations for two master variables $\psi^{\bullet}_{\ell m}$ for each spherical harmonic constructed from the components of the metric perturbation, the Regge-Wheeler (RW) and Zerilli equations [53, 54]
@@
where $\bullet=\{\mathrm{odd},\mathrm{even}\}$. The respective potentials are given by
@@
where the wave is purely ingoing at the horizon, $\psi\sim e^{-i\omega r_{\star}}$ as $r\to 2M$ ($r_{\star}\to$-$\infty$). We define the reflectivity of the BH as
@@
We implemented two methods to compute reflectivity coefficients across frequencies: one leveraging the Black Hole Perturbation Toolkit implementation of the Mano-Suzuki-Takasugi (MST) method [71, 72], and another via direct RW equation integration. Both methods agree to subpercent accuracy up to $\ell=30$, while the MST method enables reliable computations up to $\ell_{\rm max}=70$ for $\omega M\in[0.01,14]$. The reflectivity behavior is shown in Fig. 3. Notice that the reflectivity in the even parity sector is equal in magnitude to the odd sector (as required by isospectrality), and with a phase shift that can be computed through
@@
The well-defined helicity modes of the GW strain $h_{(\pm 2)}$ can be decomposed in spin weighted spherical harmonics centered around the lens
@@
$\quad\times\Bigl{[}(-1)^{\ell}e^{-i\omega r_{\star}}-e^{i\omega r_{\star}}\Bigr{]}+\mathrm{c.m.}\,$,
@@
This solution corresponds to the incoming spin-2 wave impinging the BH and passing through the system unaffected. We account for the strong field of the BH by rescaling the outgoing field with the BH reflectivity factors. In order to do so, we first subtract the incoming waveform, and then rescale the outgoing component as indicated below. We will later recover the ingoing piece, which leads to the component of the waveform that is directly propagated to the observer. Hence, we assume that the direct component does not interact with the potential of the lens. All the interaction is captured by the transmitted component, which is captured by
@@
$\quad\times\Bigl{[}1+(-1)^{\ell}\mathcal{R}^{\rm even}_{\ell m}\Bigr{]}e^{i\omega r_{\star}}+\mathrm{c.m.}\,$,
@@
$\tilde{H}$ ${}^{(\pm 2)}=\mathcal{F}^{\pm}H^{(\mp 2)}/\omega\,$, (28)
@@
$f$ ${}^{\pm}_{\ell m}=\Bigl{[}1+(-1)^{\ell}\mathcal{R}^{\rm even}_{\ell m}\Bigr{]}\mp\frac{m}{|m|}\Bigl{[}1+(-1)^{\ell}\mathcal{R}^{\rm odd}_{\ell m}\Bigr{]}\,$,
@@
For each $\ell$ mode, there are two contributions, coming from $m=\pm s$, with $s=\pm 2$ in this case. At large $\ell$, the even and odd reflectivities become nearly identical, i.e., $\mathcal{R}_{\ell m}^{\rm even}-\mathcal{R}_{\ell m}^{\rm odd}\xrightarrow{\ell\to\infty}0$. As a result, the partial sums for $m=s$, which depend on the difference between these reflectivities, exhibit rapid convergence. In contrast, for m=sm=-s the partial sums do not converge. This is a known phenomenon in the multipolar expansion of the scattering of plane waves [75, 40], and convergence is only achieved within some finite radius of the BH for a given value of $\ell_{\text{max}}$ [76]. Nevertheless, while the sum does not converge in the conventional sense, it possesses a well-defined Cesàro sum  [^3] The Cesàro sum, in its basic form, is obtained by averaging the sequence of partial sums and taking the limit. More details are given in the Appendix A., which we have verified numerically. In the Appendix A, we provide extensive numerical evidence of the sum’s effective convergence, as well as the errors introduced when truncating at finite $\ell$. We note that achieving convergence requires summing over a larger range of $\ell$ modes when the observer-lens angle $\theta_{L}$ is small. This imposes a practical limitation on probing smaller impact parameters, where richer strong-field effects will be observed.
@@ III.3 Waveform at the observer’s location
where $r_{\rm OS},\Omega_{\rm OS}$ are, respectively, the distance and the solid angle connecting the source and the observer’s locations. We note that this formulation assumes no interaction between the direct component of the emitted gravitational wave and the potential of the lens, serving as the primary simplifying assumption of this work. A more sophisticated treatment of the direct piece must be required in order to recover the geometric optics results of the PL model in the high-frequency limit. We focus on cases where the lens and source are relatively close compared to the observer, $r_{\rm LS}\ll r_{\rm OS}\sim r_{\rm OL}$. Under this approximation $\Omega_{\rm OS}\simeq\Omega_{\rm LS}$ (with equality holding in the case where source, lens, and observer are aligned). This allows us to express the total waveform compactly as
@@
The total waveform is $h=h_{+}\mathcal{A}_{+}+h_{\times}\mathcal{A}_{\times}$, where $\mathcal{A}_{+/\times}$ are the detector’s antenna response functions of the ++ and $\times$ polarizations. Including the effects of lensing, this becomes
@@
Figure 4: Absolute value of the amplification factor relative to the nonmagnified case, for the polarization preserving contribution, $|\mathcal{F}-1|^{2}$ (blue), and of the total polarization-mixing amplification $|\mathcal{G}|^{2}$ (green), as a function of the dimensionless frequency $\omega M$. The lens is located at $r_{\rm SL}=100M$, and the source has angular coordinates $(\theta_{\rm OS},\phi_{\rm OS})=(\pi/6,\pi/6)$ with respect to the observer. In red, we give $|\mathcal{F}-1|^{2}$ for the case of a PL, corresponding to an impact parameter $y=2.69$. The dimensionless frequency $\varpi$ is related to $\omega M$ via Eq. (2).
In Fig. 4, we additionally show the square of $\mathcal{F}-1$ computed using the PL approximation, for a situation matching the strong field calculation. We emphasize two main differences: (i) the strong field calculation outputs larger magnifications, relative to the PL approximation (for the polarization mixing contribution, due to our choice of the angle $\phi$), and (ii) at high frequencies, the amplification factor $|\mathcal{F}|\to 1$ and $\mathcal{G}\to 0$, since absorption by the BH becomes important. This effect is unaccounted for in the PL approximation. Nevertheless, it is remarkable that a simple approximation such as the PL can reproduce accurately several of the features observed in the complete, strong field calculation.
@@ IV Waveform comparison
Once we have constructed the amplification factors, we can directly compute the lensed waveform and analyze its most prominent features. Our amplification factor can be evaluated efficiently across a range of $\omega M$, and thus we capture key lensing effects in the strong gravitational regime throughout the entire inspiral merger ringdown. We will examine both the strong-field regime and the weak-field, PL approximation. For clarity, we will focus on a specific gravitational waveform resulting from the merger of spinless, equal mass BHs. This allows us to isolate the main contributions from strong-field lensing compared to the weak-field approximation. We will explore the lensed waveform in both the frequency and time domains.